STRUCTURAL ANALYSIS AND DIPOLE MOMENT DETERMINATION OF THE GAS-PHASE TRIMETHYLAMINE SULFUR TRIOXIDE COMPLEX

\underline{D. L. FIACCO}, M. E. OTT, and K. R. LEOPOLD, Department of Chemistry, University of Minnesota, Minneapolis, MN 55455; A. TORO, Pontifical Catholic University of Puerto Rico, Ponce, PR 00731-6382.

The structure and dipole moment of the Lewis acid-base adduct (CH_{3})_{3}N-SO_{3} have been determined by microwave spectroscopy. Analysis yields an N-S bond length of 1.914(33) Å and an NSO angle of 100.0(4)^{∘}. These data indicate that the formation of the N-S dative bond is nearly (but not entirely) complete in the gas phase. Comparison with previous work on the closely related system H_{3}N-SO_{3} indicates that the increased basicity of the (CH_{3})_{3}N relative to that of H_{3}N is effective at driving the dative bond further towards completion. Analysis of nuclear hyperfine structure indicates that about 0.6 electrons are transferred from the nitrogen to the SO_{3} upon formation of the complex. This is significantly larger than that observed in H_{3}N-SO_{3}. The dipole moment, obtained from Stark effect measurements, of 7.111(7) D, arises from a combination of charge transfer, out-of-plane distortion of the SO_{3} and the intrinsic moment of (CH_{3})_{3}N.